OpenStudy (anonymous):

anybody know how to solve log^3(5)-log^3(x+8)=2

myininaya (myininaya):

thats base 3 right?

myininaya (myininaya):

since they have the same base you write log_3(5/(x+8))=2

myininaya (myininaya):

and then write base 3 on both sides to obtain 5/(x+8)=3^2

OpenStudy (anonymous):

okay and then do you just solve for x from that equation? sweet. Thank you anybody know

OpenStudy (anonymous):

how do you simplify (6p/2p+5)-(5/p+6)